If 19.22 grams of aluminum is heated from 15 oC to 50 oC how much energy was required?

sAl = 0.897 J/(gK)


1
Expert's answer
2021-03-09T04:59:55-0500

Solution:

q = m × C × ΔT

q = m × C × (Tf - Ti),

where:

q = amount of heat energy gained or lost by substance

m = mass of sample

C = specific heat capacity (J oC-1 g-1)

Tf = final temperature

Ti = initial temperature

 

Thus:

q = (19.22 g) × (0.897 J oC-1 g-1) × (50 - 15)oC = 603.4 J

q = 603.4 J (or 0.6 kJ)


Answer: 603.4 Joules of energy was required.